OpenStudy (anonymous):

Simplify the expression 1 over 4 to the negative fourth power.

OpenStudy (anonymous):

\[\frac{ 1 }{ 4^{-4} }\] is that the question

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[x^a= \frac{ 1 }{ x^{-a} }\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

so it would just be 4^4 as in 256?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!